OpenStudy (anonymous):

How much current does a 48 W device draw when 120 V is placed across it? A. 4.0 * 10^-1 A B. 2.5 * 10^0 A C. 4.2 * 10^2 A D. 3.6 * 10^4 A

OpenStudy (michele_laino):

power is equal to voltage times current, so we have: I=W/V=48/120=...amperes

OpenStudy (anonymous):

0.4!

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

so choice A is our solution?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

thanks!!:D

OpenStudy (michele_laino):

:)